a ball is thrown vertically upwards from the foot of a tower.it crosses the top of the tower twice after an interval of 4s and reaches the foot of the tower 8s after it was thrown.what is the height of the tower??

Grade:12

2 Answers

Arun
4 years ago
Dear Student
 
time taken to go upto U and then come back to the ground required 8 sec.
So 4 s up then 4 s in coming down.
the time taken between climbing from T to U and then coming to T again(crossing T twice) =4 s
so time taken to go upto T from the ground=(8 - 4)/2=2s
now if we consider the journey from ground to U:
0=u-4x10 ....... putting in 1st eq.
u=40 m/s
now tower height:H=
H=40x2-1/2 x10x2x2.....putting in second eq
H=60 m
 
Vikas TU
4 years ago